### v1

# LoRA-IR: Taming Low-Rank Experts for Efficient All-in-One Image Restoration

Abstract
Prompt-based all-in-one image restoration (IR) frameworks have achieved remarkable performance by incorporating degradation-specific information into prompt modules. Nevertheless, handling the complex and diverse degradations encountered in real-world scenarios remains a significant challenge. To address this challenge, we propose LoRA-IR, a flexible framework that dynamically leverages compact low-rank experts to facilitate efficient all-in-one image restoration. Specifically, LoRA-IR consists of two training stages: degradation-guided pre-training and parameter-efficient fine-tuning. In the pre-training stage, we enhance the pre-trained CLIP model by introducing a simple mechanism that scales it to higher resolutions, allowing us to extract robust degradation representations that adaptively guide the IR network. In the fine-tuning stage, we refine the pre-trained IR network using low-rank adaptation (LoRA). Built upon a Mixture-of-Experts (MoE) architecture, LoRA-IR dynamically integrates multiple low-rank restoration experts through a degradation-guided router. This dynamic integration mechanism significantly enhances our model’s adaptability to diverse and unknown degradations in complex real-world scenarios. Extensive experiments demonstrate that LoRA-IR achieves state-of-the-art performance across 14 image restoration tasks and 29 benchmarks. Code and pre-trained models will be available at: https://github.com/shallowdream204/LoRA-IR.
Corresponding author.
1 Introduction
Image restoration (IR) is a fundamental task in computer vision, aiming to recover high-quality (HQ) images from degraded low-quality (LQ) inputs. In recent years, significant progress has been achieved with specialized restoration networks targeting specific degradations such as adverse weather removal [15, 7], denoising [83], deblurring [58], and low-light enhancement [70]. However, in practical applications like autonomous driving and outdoor surveillance [40, 96], images are often simultaneously affected by multiple complex degradations, including haze, rain, snow, low-light conditions, and motion blur. These complex degradations not only degrade image quality but also severely impair the performance of downstream vision tasks, posing significant challenges to the safety and reliability of systems. Existing specialized models designed for single-task restoration struggle to generalize effectively in such unpredictable and variable environments.
Figure 1: PSNR comparison with state-of-the-art all-in-one methods across 8 tasks (Tab. 4 and Tab. 9).
Figure 2: Conceptual comparison of all-in-one frameworks. (a) Multi-Encoder Structures: Use multiple encoders to extract features, but redundancy reduces computational efficiency. (b) Prompt-Based Methods: Employ lightweight prompts for degradation-specific features, improving efficiency. However, static network structures limit their ability to handle unknown complex degradations. (c) Our Proposed Framework: Dynamically and sparsely combines low-rank restoration experts. This design maintains computational efficiency while adapting flexibly to diverse degradation types, enhancing real-world performance.
To overcome the limitations of specialized models, there is increasing interest in developing all-in-one frameworks capable of handling diverse degradations. Early approaches, such as multi-encoder architectures [24] (Fig. 2 (a)), employ separate encoders for different degradation types. While effective in handling multiple degradations, their redundant structures lead to a large number of parameters, hindering scalability and efficiency. More recent methods adopt prompt-based frameworks [48, 33, 36, 2] (Fig. 2 (b)), encoding degradation-specific information into lightweight prompts to guide a shared network. However, relying solely on lightweight prompts and a static shared network may not fully capture the fine-grained details and specific patterns associated with different degradations, leading to suboptimal restoration results. Furthermore, potential correlations and shared features among different degradations—such as common patterns in adverse weather conditions [95, 80]—are not extensively leveraged. Exploring these correlations could be key to enhancing model adaptability and effectiveness in complex real-world scenarios.
In this work, we propose LoRA-IR, a flexible framework for efficient all-in-one image restoration (Fig. 2 (c)). Motivated by the success of Low-Rank Adaptation (LoRA) [14] in parameter-efficient fine-tuning, this paper explores the use of diverse low-rank experts to model degradation characteristics and correlations efficiently. LoRA-IR consists of two stages of training, both guided by a proposed degradation-guided router (DG-Router). DG-Router is based on the powerful vision-language model CLIP [51], which has demonstrated strong representational capabilities across a wide range of high-level vision tasks [29, 90]. However, when applied to low-level tasks, its limited input resolution inevitably leads to suboptimal performance when handling high-resolution LQ images. To this end, we introduce a simple yet effective method for scaling CLIP to high resolution. Our approach involves downsampling the image and applying a sliding window technique to capture both global and local detail representations, which are subsequently fused using lightweight MLPs. With minimal training parameters and a short training time, DG-Router provides robust degradation representations and probabilistic guidance for the training of LoRA-IR.
In the first stage, we use the degradation representations provided by the DG-Router to guide the pre-training of the IR network. The degradation representations dynamically modulate features within the IR network through the proposed degradation-guided adaptive modulator (DAM). In the second stage, we fine-tune the IR network obtained from the first stage using LoRA. Based on the Mixture-of-Expert (MoE) [53] structure, we construct a set of low-rank restoration experts. Leveraging the probabilistic guidance of the DG-Router, we sparsely select different LoRA experts to dynamically adjust the IR network. Different experts enhance the network’s ability to capture degradation-specific knowledge, while their collaboration equips the network with the capability to learn correlations between various degradations. The dynamic network structure enables LoRA-IR to adapt to diverse degradations and improves its generalization capabilities. As shown in Fig. 1, LoRA-IR outperforms all compared state-of-the-art all-in-one methods and demonstrates favorable generalizability in handling complex real-world scenarios.
The main contributions can be summarized as follows:
  • We propose LoRA-IR, a simple yet powerful baseline for all-in-one restoration. LoRA-IR employs a novel mixture of low-rank expert structures, boosting architectural flexibility while maintaining computational efficiency.
  • We propose a CLIP-based degradation-guided router (DG-Router) to extract robust degradation representations. DG-Router requires minimal training parameters and time, offering valuable guidance for LoRA-IR.
  • Extensive experiments across 14 image restoration tasks and 29 benchmarks validate the SOTA performance of LoRA-IR. Notably, LoRA-IR exhibits strong generalizability to real-world scenarios, including training-unseen tasks and mixed-degradation removal.
2 Related Work
2.1 Image Restoration
Image restoration for known degradations has been extensively studied [27, 77, 78, 65, 4, 82, 26, 71, 9]. Recently, there has been significant interest in all-in-one frameworks within the community [81, 95, 60]. AirNet [22] is the pioneering work in all-in-one IR, utilizing contrastive learning to capture degradation information. Recent SOTA methods are mostly based on prompt learning, using lightweight prompts to encode degradation information. PromptIR [48] proposes a plug-and-play prompt module to guide the restoration process. DA-CLIP [36] utilizes a prompt learning module to incorporate degradation embeddings. MPerceiver [2] introduces a multi-modal prompt learning approach to harness Stable Diffusion priors. Despite achieving promising results, most existing methods use fixed network architectures, which may limit their ability to cover complex real-world scenarios.
2.2 Vision-Language Models
Vision-language models (VLMs) have demonstrated strong performance across various multi-modal and vision-only tasks [51, 8, 29]. Among them, CLIP, as a powerful VLM, has demonstrated impressive zero-shot and few-shot capabilities across various high-level vision tasks [67, 93, 66]. However, in low-level vision tasks, CLIP’s capabilities have been relatively less explored. DA-CLIP [36] is the first to introduce CLIP into all-in-one IR. It adopts a ControlNet-style [86] structure, using contrastive learning with image-text pairs to fine-tune CLIP. In this work, we focus on exploring CLIP’s visual representation capabilities to achieve efficient degradation representation. Compared to DA-CLIP (Tab. 7), our proposed DG-Router requires 64×64$\times$64 × fewer learning parameters and 4×4$\times$4 × less training time, while achieving superior performance.
2.3 Parameter-efficient Fine-tuning
With the rise of large foundational models [51, 18, 1] in modern deep learning, the community has increasingly shifted its focus towards parameter-efficient fine-tuning (PEFT) methods for effective model adaptation. Among these, prompt learning [20, 90] and Low-Rank Adaptation (LoRA) [14] are two prominent and widely used PEFT methods. As discussed above, prompt learning has been widely applied in low-level vision tasks. LoRA posits that the weight changes during model adaptation follow a low-rank structure and incorporates trainable rank decomposition matrices into the pre-trained model. Specifically, the change matrix is re-parameterized into the product of two low-rank matrices: $W=W_{0}+\Delta W=W_{0}+sBA,$ where $W_{0}$ represents pre-trained weight matrix, $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are low-rank matrices, $s=\frac{\alpha}{r}$ is the scaling factor. In this work, we first introduce LoRA into the all-in-one frameworks to facilitate efficient image restoration.
Figure 3: Visualization of images output by the CLIP processor (top row from GoPro [44] and bottom row from LIVE1 [54]), which reveals significant loss of degradation information after processing. Please zoom in for a better view.
Figure 4: Overall of the proposed LoRA-IR, which includes (a) Degradation-guided router (DG-Router), (b) Pre-training image restoration network with robust degradation embedding, (c) Fine-tuning image restoration network with low-rank restoration experts, and (d) Degradation-guided adaptative modulator (DAM).
3 Method
As shown in Fig. 4, the image restoration network is based on the commonly used U-Net [65, 78, 4] structure, consisting of stacked encoder, middle, and decoder blocks. Our LoRA-IR consists of two training stages: degradation-guided pre-training and parameter-efficient fine-tuning, both directed by the proposed degradation-guided router (DG-Router). We first introduce the CLIP-based DG-Router in Sec. 3.1, which is used to extract robust degradation representations and provide probabilistic estimates to guide the training of LoRA-IR. Then we detail the pre-training process of LoRA-IR in Sec. 3.2. Finally, we describe the fine-tuning process in Sec. 3.3.
3.1 Degradation-guided Router
As shown in Fig. 4 (a), DG-Router uses a pre-trained CLIP image encoder to extract rich features from LQ images. The pre-trained CLIP image encoder typically limits input images to a smaller resolution (e.g., $224\times 224$). When handling higher-resolution images, a common approach [36] is to downsample the image to the resolution supported by CLIP using the processor. While this may have minimal impact on perception-based high-level classification tasks, significant downsampling can potentially lead to the loss of critical degradation information in pixel-level regression tasks like image restoration. Fig. 3 illustrates the results after the CLIP processor has processed the LQ images. It can be seen that significant downsampling leads to a substantial loss of degradation information in the LQ images, making it challenging to effectively extract degradation representations from the CLIP output features.
To address this issue, we propose a simple yet effective mechanism for scaling up the input resolution. For input LQ image $I_{LQ}\in\mathbb{R}^{H\times W\times 3}$, we use a sliding window to partition the image into small local patches $I_{slide}\in\mathbb{R}^{M\times H_{c}\times W_{c}\times 3}$, where $M$ is the number of patches, $H_{c}\times W_{c}$ is the CLIP supported resolution. Both $I_{slide}$ and down-sampled image $I_{down}\in\mathbb{R}^{H_{c}\times W_{c}\times 3}$ are fed into the image encoder simultaneously, obtaining output features $e^{slide}\in\mathbb{R}^{M\times C_{clip}}$ and $e^{down}\in\mathbb{R}^{C_{clip}}$. As depicted in Fig. 4 (a), after pooling $e^{slide}$, we concatenate them and pass them into a two-layer MLP to obtain the CLIP extracted degradation embedding $e^{clip}$, which can be formulated as
$\begin{split}[&e^{down},e^{slide}]=\mathrm{CLIP}([I_{down},I_{slide}]),\\ &e^{clip}=\mathrm{MLP}(\mathrm{Concat}(e^{down},\mathrm{Pooling}(e^{slide}))).\end{split}$ (1)
After feeding $e^{clip}$ into the classification head, we ultimately obtain the degradation prediction probability $w\in\mathbb{R}^{n}$, where $n$ is the number of degradation types. Without bells and whistles, the DG-Router is optimized using standard cross-entropy loss, with the only parameters being the classification head and the two-layer MLP. Once training is complete, all parameters of the DG-Router are frozen and no longer updated.
3.2 Degradation-guided Pre-training
In the pre-training stage (Fig. 4 (b)), We dynamically modulate the restoration network using the degradation representations $e^{clip}$ extracted by the DG-Router. We propose a degradation-guided adaptive modulator (DAM) to modulate the features of the restoration network. As shown in Fig. 4 (d), we first use a two-layer MLP projector to transform $e^{clip}$ into a degradation embedding $e^{d}$ in the feature space of the IR network. DAM adopts a structure similar to the channel attention block [88], modulating degradation information along the channel dimension, which can be formulated as
$\begin{split}e^{d}&=\mathrm{MLP_{shared}}(e^{clip}),\\ x_{out}&=\mathrm{LN}(x_{in})\odot\mathrm{Sigmoid}(\mathrm{MLP}(e^{d}))+x_{in},\end{split}$ (2)
where $\odot$ denotes the element-wise multiplication, $\mathrm{MLP_{shared}}$ denotes the MLP projector shared across different blocks, $\mathrm{LN}$ denotes LayerNorm, $x_{in}$ is the original feature in the IR network, and $x_{out}$ is the feature after modulation. Through the modulation of DAM, the robust degradation representations from the DG-Router can effectively enhance the degradation-specific knowledge of the IR network during pre-training.
3.3 Parameter-efficient Fine-tuning
In the fine-tuning stage, we aim to utilize the Low-Rank Adaptation (LoRA) technique to model degradation characteristics and correlations efficiently, enhancing the model’s ability to handle training-seen degradations and improving its robustness for real-world training-unseen degradations.
As shown in Fig. 4 (c), built upon the Mixture-of-Expert (MoE) architecture, we establish a collection composed of low-rank restoration experts. We have a total of $n$ low-rank experts $\{E_{1},E_{2},\cdots,E_{n}\}$, each expert is a learnable lightweight LoRA weight of the pre-trained restoration network from the first stage, specialized in handling a specific type of degradation.
For a given input LQ image, DG-Router can predict its degradation probability $w\in\mathbb{R}^{n}$. $w$ is used as the score of different experts to select the appropriate expert for the restoration process. We sparsely select the top-$k$ highest-scoring experts as the most relevant ones and achieve the final restoration result through their dynamic collaboration, which can be formulated as
$x_{out}=PreMod(x_{in})+\sum_{i=1}^{k}w_{\varphi(i)}^{\prime}E_{\varphi(i)}(x_{in}),$ (3)
where $PreMod$ denotes the pre-trained module in stage1, $\varphi(i)$ denotes the index of the $i$-th selected expert, $w^{\prime}\in\mathbb{R}^{n}$ represents the result of reapplying softmax normalization to the scores of the selected top-$k$ experts (the weights of the experts that were not selected are set to 00).
Note that the sparse selection mechanism in Eq. (3) grants LoRA-IR a flexible network structure, enhancing its ability to represent degradation-specific knowledge. The dynamic combination mechanism, on the other hand, facilitates collaboration among different restoration experts, effectively modeling the commonalities and correlations between various degradations. The design of the low-rank experts ensures the efficiency of LoRA-IR, enabling it to achieve high-performance all-in-one image restoration in a computationally efficient manner.
4 Experiments
Table 1: [Setting I] Quantitative comparisons for 4-task adverse weather removal. The top halves of the tables show task-specific results, while the bottom halves present all-in-one evaluations. LoRA-IR surpasses recent SOTA all-in-one techniques, including MPerceiver [2] and Histoformer [56], across all evaluated datasets and metrics.
(a) Image Desnowing
Snow100K-S [34] Snow100K-L [34]
PSNR SSIM PSNR SSIM
SPANet [63] 29.92 0.8260 23.70 0.7930
JSTASR [5] 31.40 0.9012 25.32 0.8076
RESCAN [25] 31.51 0.9032 26.08 0.8108
DesnowNet [34] 32.33 0.9500 27.17 0.8983
DDMSNet [84] 34.34 0.9445 28.85 0.8772
NAFNet [4] 34.79 0.9497 30.06 0.9017
Restormer [78] 36.01 0.9579 30.36 0.9068
All-in-One [24] - - 28.33 0.8820
TransWeather [60] 32.51 0.9341 29.31 0.8879
Chen et al. [6] 34.42 0.9469 30.22 0.9071
WGWSNet [95] 34.31 0.9460 30.16 0.9007
WeatherDiff64 [46] 35.83 0.9566 30.09 0.9041
WeatherDiff128 [46] 35.02 0.9516 29.58 0.8941
AWRCP [75] 36.92 0.9652 31.92 0.9341
MPerceiver [2] 36.23 0.9571 31.02 0.9164
Histoformer [56] 37.41 0.9656 32.16 0.9261
LoRA-IR (Ours) 37.89 0.9683 32.28 0.9296
(b) Deraining & Dehazing
Outdoor-Rain [23]
PSNR SSIM
CycleGAN [94] 17.62 0.6560
pix2pix [16] 19.09 0.7100
HRGAN [23] 21.56 0.8550
PCNet [17] 26.19 0.9015
MPRNet[77] 28.03 0.9192
NAFNet [4] 29.59 0.9027
Restormer [78] 30.03 0.9215
All-in-One [24] 24.71 0.8980
TransWeather [60] 28.83 0.9000
[6] 29.27 0.9147
WGWSNet [95] 29.32 0.9207
WeatherDiff64 [46] 29.64 0.9312
WeatherDiff128 [46] 29.72 0.9216
AWRCP [75] 31.39 0.9329
MPerceiver [2] 31.25 0.9246
Histoformer [56] 32.08 0.9389
LoRA-IR (Ours) 32.62 0.9447
(c) Raindrop Removal
RainDrop [49]
PSNR SSIM
pix2pix [16] 28.02 0.8547
DuRN [32] 31.24 0.9259
RaindropAttn [50] 31.44 0.9263
AttentiveGAN [49] 31.59 0.9170
IDT [72] 31.87 0.9313
MAXIM [59] 31.87 0.9352
Restormer [78] 32.18 0.9408
All-in-One [24] 31.12 0.9268
TransWeather [60] 30.17 0.9157
[6] 31.81 0.9309
WGWSNet [95] 32.38 0.9378
WeatherDiff64 [46] 30.71 0.9312
WeatherDiff128 [46] 29.66 0.9225
AWRCP [75] 31.93 0.9314
MPerceiver [2] 33.21 0.9294
Histoformer [56] 33.06 0.9441
LoRA-IR (Ours) 33.39 0.9489
Table 2: [Setting III] Quantitative comparison with all-in-one models for 3-task image restoration.
Method Dehazing Deraining Denoising on BSD68 dataset Average
on SOTS on Rain100L $\sigma=15$ $\sigma=25$ $\sigma=50$
BRDNet [57] 23.23 / 0.895 27.42 / 0.895 32.26 / 0.898 29.76 / 0.836 26.34 / 0.836 27.80 / 0.843
LPNet [12] 20.84 / 0.828 24.88 / 0.784 26.47 / 0.778 24.77 / 0.748 21.26 / 0.552 23.64 / 0.738
FDGAN [10] 24.71 / 0.924 29.89 / 0.933 30.25 / 0.910 28.81 / 0.868 26.43 / 0.776 28.02 / 0.883
MPRNet [77] 25.28 / 0.954 33.57 / 0.954 33.54 / 0.927 30.89 / 0.880 27.56 / 0.779 30.17 / 0.899
DL [11] 26.92 / 0.391 32.62 / 0.931 33.05 / 0.914 30.41 / 0.861 26.90 / 0.740 29.98 / 0.875
AirNet [22] 27.94 / 0.962 34.90 / 0.967 33.92 / 0.933 31.26 / 0.888 28.00 / 0.797 31.20 / 0.910
PromptIR [48] 30.58 / 0.974 36.37 / 0.972 33.98 / 0.933 31.31 / 0.888 28.06 / 0.799 32.06 / 0.913
LoRA-IR (Ours) 30.68 / 0.961 37.75 / 0.979 34.06 / 0.935 31.42 / 0.891 28.18 / 0.803 32.42 / 0.914
Table 3: [Setting II] Quantitative comparisons for 3-task real-world adverse weather removal on WeatherStream [80].
Method Rain Haze Snow Average
MPRNet [77] 21.50 21.73 20.74 21.32
NAFNet [4] 23.01 22.20 22.11 22.44
Uformer [65] 22.25 18.81 20.94 20.67
Restormer [78] 23.67 22.90 22.51 22.86
GRL [26] 23.75 22.88 22.59 23.07
AirNet [22] 22.52 21.56 21.44 21.84
TUM [6] 23.22 22.38 22.25 22.62
Transweather [60] 22.21 22.55 21.79 22.18
WGWS [95] 23.80 22.78 22.72 23.10
LDR [73] 24.42 23.11 23.12 23.55
LoRA-IR (Ours) 25.22 24.39 23.31 24.31
4.1 Experimental Setup
Settings. To comprehensively evaluate our method, we conduct experiments in five different settings following previous works: (I) 4-task adverse weather removal [56], including desnowing, deraining, dehazing, and raindrop removal; (II) 3-task real-world adverse weather removal [73], including deraining, dehazing, and desnowing; (III) 3-task image restoration [22], including deraining, dehazing, and denoising; (IV) 5-task image restoration [89], including deraining, low-light enhancement, desnowing, dehazing, and deblurring; (V) 10-task image restoration [36], including deblurring, dehazing, JPEG artifact removal, low-light enhancement, denoising, raindrop removal, deraining, shadow removal, desnowing, and inpainting. For each setting, we train a unified model to handle multiple types of degradations.
Datasets and Metrics. For Setting I, we use the AllWeather [60, 56] dataset to evaluate our method. For Setting II, we use the WeatherStream [80] dataset to evaluate the model’s performance in real-world scenarios. For Setting III, we use RESIDE [21] for dehazing, WED [39] and BSD [41] for denoising, Rain100L [74] for deraining. For Setting IV, we use a merged dataset [78, 89] for deraining, LOL [68], DCIE [19], MEF [38], and NPE [62] for low-light enhancement, Snow100K [34] for desnowing, RESIDE [21] for dehazing, GoPro [44], HIDE [55], RealBlur [52] for deblurring. For Setting V, we use the same dataset as [36]. We provide a detailed introduction to the datasets used for training and testing in Appendix.
As for evaluation metrics, we adopt PSNR and SSIM as the distortion metrics, LPIPS [87] and FID [13] as perceptual metrics. For benchmarks that do not include ground truth images, we use NIQE [42], LOE [61] and IL-NIQE [85] as no-reference metrics.
Implementation Details. For the training of DG-Router, we use the Adam optimizer with a batch-size of $64\times n$, where $n$ is the number of tasks. The whole training takes 20 minutes with a fixed learning rate of $2e^{-4}$ using 8 NVIDIA A100 GPUs. Our LoRA-IR follows a two-stage training process, namely pre-training and fine-tuning. For both stages, we use the AdamW optimizer with a batch-size of 64. Following [36, 89], the training patch size is set to $256$ to ensure fair comparisons. Random cropping, flipping, and rotation are used as data augmentation techniques. For the image restoration network structure, all basic blocks in Fig. 4 are the simple convolutional NAFBlocks [4], forming a simple all-in-one CNN baseline. More specific details for different settings provided in Appendix.
4.2 Comparison with State-of-the-Art Methods
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Blurry Image Input PromptIR [48] DiffUIR [89] LoRA-IR (Ours) GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LQ Image Input PromptIR [48] DiffUIR [89] LoRA-IR (Ours) GT
Figure 5: [Setting IV] Visual results on HIDE [55] for training-seen tasks generalization evaluation (top row) and TOLED [92] for training-unseen tasks generalization evaluation (bottom row). Zoom in for a better view.
Table 4: [Setting IV] Comparison with state-of-the-art task-specific and all-in-one methods for 5-task image restoration.
Method Deraining $(5sets)$ Enhancement Desnowing $(2sets)$ Dehazing Deblurring
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Task-Specific
SwinIR [27] - - 17.81 0.723 - - 21.50 0.891 24.52 0.773
MIRNetV2 [79] - - 24.74 0.851 - - 24.03 0.927 26.30 0.799
IR-SDE [35] - - 20.45 0.787 - - - - 30.70 0.901
WeatherDiff [46] - - - - 33.51 0.939 - - - -
RDDM [30] 30.74 0.903 23.22 0.899 32.55 0.927 30.78 0.953 29.53 0.876
All-in-One
Restormer [78] 27.10 0.843 17.63 0.542 28.61 0.876 22.79 0.706 26.36 0.814
AirNet [22] 24.87 0.773 14.83 0.767 27.63 0.860 25.47 0.923 26.92 0.811
Painter [64] 29.49 0.868 22.40 0.872 - - - - - -
IDR [81] - - 21.34 0.826 - - 25.24 0.943 27.87 0.846
ProRes [37] 30.67 0.891 22.73 0.877 - - - - 27.53 0.851
PromptIR [48] 29.56 0.888 22.89 0.847 31.98 0.924 32.02 0.952 27.21 0.817
DACLIP-UIR [36] 28.96 0.853 24.17 0.882 30.80 0.888 31.39 0.983 25.39 0.805
DiffUIR-L [89] 31.03 0.904 25.12 0.907 32.65 0.927 32.94 0.956 29.17 0.864
LoRA-IR (Ours) 32.35 0.924 26.42 0.926 34.16 0.941 35.74 0.986 30.09 0.914
Table 5: [Setting IV] Comparison on real-world benchmarks for training-seen tasks generalization evaluation. Best and second best performance of all-in-one approaches are marked in bold and underlineded, respectively.
Method Deraining Enhancement Desnowing Deblurring
[…]
Task-Specific
WeatherDiff [46] - - - - 2.96 21.976 - -
CLIP-LIT [28] - - 3.70 232.48 - - - -
RDDM [30] 3.34 41.80 3.57 202.18 2.76 22.261 30.74 0.894
Restormer [78] 3.50 30.32 3.80 351.61 - - 32.12 0.926
All-in-One
AirNet [22] 3.55 145.3 3.45 598.13 2.75 21.638 16.78 0.628
PromptIR [48] 3.52 28.53 3.31 255.13 2.79 23.000 22.48 0.770
DACLIP-UIR [36] 3.52 42.03 3.56 218.27 2.72 21.498 17.51 0.667
DiffUIR [89] 3.38 24.82 3.14 193.40 2.74 22.426 30.63 0.890
LoRA-IR (Ours) 3.47 67.53 3.28 93.32 2.70 22.010 30.80 0.907
Table 6: [Setting IV] Comparison on TOLED and POLED datasets [92] for training-unseen tasks generalization evaluation (under-display camera image restoration).
Method TOLED [92] POLED [92]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NAFNet [4] 26.89 0.774 0.346 10.83 0.416 0.794
HINet [3] 13.84 0.559 0.448 11.52 0.436 0.831
MPRNet [77] 24.69 0.707 0.347 8.34 0.365 0.798
DGUNet [43] 19.67 0.627 0.384 8.88 0.391 0.810
MIRNetV2 [79] 21.86 0.620 0.408 10.27 0.425 0.722
SwinIR [27] 17.72 0.661 0.419 6.89 0.301 0.852
RDDM [30] 23.48 0.639 0.383 15.58 0.398 0.544
Restormer [78] 20.98 0.632 0.360 9.04 0.399 0.742
DL [11] 21.23 0.656 0.434 13.92 0.449 0.756
Transweather [60] 25.02 0.718 0.356 10.46 0.422 0.760
TAPE [31] 17.61 0.583 0.520 7.90 0.219 0.799
AirNet [22] 14.58 0.609 0.445 7.53 0.350 0.820
IDR [81] 27.91 0.795 0.312 16.71 0.497 0.716
PromptIR [48] 16.70 0.688 0.422 13.16 0.583 0.619
DACLIP-UIR [36] 15.74 0.606 0.472 14.91 0.475 0.739
DiffUIR-L [89] 29.55 0.887 0.281 15.62 0.424 0.505
LoRA-IR (Ours) 28.68 0.876 0.279 17.02 0.700 0.600
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Low-light&Blurry Input DACLIP-UIR [36] DiffUIR [89] LoRA-IR (Ours) GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Rainy&Hazy Input DACLIP-UIR [36] DiffUIR [89] LoRA-IR (Ours) GT
Figure 6: Visual comparisons on challenging mixed-degradation benchmarks. Please zoom in for a better view.
Setting I. Tab. 1 shows the comparison results of our method with task-specific methods and all-in-one methods. Compared to SOTA methods like MPerceiver [2] and Histoformer [56], our approach shows significant improvements across all benchmarks and metrics. Visual comparisons are provided in the Appendix.
Setting II. To further demonstrate the effectiveness of our method in mitigating real-world adverse weather conditions, we evaluate its performance on the WeatherStream [80] dataset. Tab. 3 presents the quantitative comparison results of PSNR with SOTA general IR as well as all-in-one IR methods. Compared to the SOTA method LDR [73], our method achieves an average PSNR improvement of 0.76 dB across the three tasks. Visual comparisons are provided in the Appendix.
Setting III. Tab. 2 presents the quantitative comparison results for 3-task image restoration. Our method surpasses PromptIR [48] by 1.38 dB in PSNR and 0.007 in SSIM on the Rain100L dataset, with an average improvement of 0.36 dB across the three tasks. Visual comparisons are provided in the Appendix.
Setting IV. Tab. 4 presents the quantitative comparison results of our method against SOTA task-specific methods and all-in-one methods across five tasks. It shows that our method outperforms the compared all-in-one methods across all tasks and even surpasses task-specific methods in the majority of tasks. For example, compared to the recent SOTA method DiffUIR [89], LoRA-IR brings a PSNR improvement ranging from 0.92 to 2.8 dB across various tasks.
To further validate the generalizability of our method for complex degradations in real-world scenarios, we evaluate from two perspectives:
(1) Generalization on Training-seen Tasks: We directly test the trained all-in-one model on real-world benchmarks that were not seen during training. As shown in Tab. 5, our method achieves the best PSNR and SSIM metrics in deblurring tasks. As discussed in  [76, 69], diffusion-based IR methods typically have an advantage in no-reference metrics like NIQE. However, our CNN-based model achieves comparable or even better performance on no-reference metrics compared to two SOTA diffusion-based methods, DACLIP-UIR [36] and DiffUIR [89]. Notably, our method shows approximately a 100-point improvement in LOE performance over DiffUIR in enhancement. Fig. 5 also shows that our method achieves more pleasing visual results.
(2) Generalization on Training-unseen Tasks: We directly test all-in-one models on the training-unseen under-display camera image restoration task. Tab. 6 shows that, compared to general IR and all-in-one methods, our method achieves either the best or second-best performance across all metrics. Fig. 5 shows that our method produces the clearest result when handling unknown degradations.
Setting V. Tab. 7 shows that, compared to DA-CLIP [36], our DG-Router requires significantly fewer (approximately 64×64$\times$64 ×) training parameters and a shorter (about 4×4$\times$4 ×) training time, while achieving more accurate degradation predictions. As shown in Tab. 8, our LoRA-IR outperforms all compared general IR and all-in-one models in both distortion and perceptual metrics, showcasing the superiority of LoRA-IR. The detailed results for each task are provided in the Appendix due to the page limit.
Mixed-degradataion Removal. Considering that images in real-world scenarios may not contain only a single type of degradation, we further evaluated different all-in-one methods on mixed-degradation benchmarks. Our experiments include three mixed-degradation benchmarks: rain&haze [23], low-light&blur [91], and blur&JPEG [45]. Tab. 9 shows that our method has a significant advantage in handling challenging mixed-degradation scenarios. We provide visual results in Fig. 6, showcasing the effectiveness of our method in handling mixed degradations.
Table 7: [Setting V] Comparison with DA-CLIP [36] on degradation prediction. Training time is evaluated using A100 GPU hours.
Method Trainable Params Training Time Blur Haze JPEG LL Noise RD Rain Shadow Snow Inpaint
DA-CLIP 94.94M 12 91.6 100 100 100 100 100 100 100 100 100
DG-Router 1.48M 2.67 100 100 100 100 100 100 100 100 100 100
Table 8: [Setting V] Quantitative comparisons for 10-task image restoration. Our LoRA-IR outperforms SOTA methods in both distortion and perceptual metrics.
Method Distortion Perceptual
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
NAFNet [4] 26.34 0.847 0.159 55.68
Restormer [78] 26.43 0.850 0.157 54.03
IR-SDE [35] 23.64 0.754 0.167 49.18
AirNet [22] 25.62 0.844 0.182 64.86
PromptIR [48] 27.14 0.859 0.147 48.26
DACLIP-UIR [36] 27.01 0.794 0.127 34.89
LoRA-IR (Ours) 28.64 0.878 0.118 34.26
Table 9: Quantitative comparison with SOTA all-in-one methods on mixed-degradtaion benchmarks. Note that none of the models are trained on mixed-degradation data.
Method Rain & Haze [23] Low-light & Blur [91] Blur & JPEG [45]
[…]
AirNet [22] 14.02 0.627 0.477 13.84 0.611 0.344 22.79 0.692 0.389
PromptIR [48] 14.75 0.634 0.454 17.61 0.681 0.317 24.98 0.710 0.403
DACLIP-UIR [36] 15.19 0.637 0.481 15.03 0.625 0.330 24.26 0.704 0.358
DiffUIR [89] 14.87 0.631 0.459 15.97 0.657 0.339 24.86 0.714 0.332
LoRA-IR (Ours) 15.41 0.642 0.445 20.59 0.719 0.305 25.05 0.716 0.359
Table 10: Ablations of LoRA-IR on the AllWeather [60] and mixed-degradation benchmarks (Mixed1: low-light&blur [91], Mixed2: blur&jpeg [45]).
Snow [34] Rain [23] Raindrop [49] Mixed1[91] Mixed2[45]
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
DG-Router 32.28 0.930 32.62 0.945 33.39 0.949 20.59 0.719 25.05 0.716
w/o high reso. 32.19 0.929 32.31 0.942 33.20 0.947 19.51 0.711 24.94 0.714
DAM 32.28 0.930 32.62 0.945 33.39 0.949 20.59 0.719 25.05 0.716
w/o DAM 32.07 0.927 32.28 0.941 33.12 0.945 18.91 0.704 24.89 0.711
AdaLN [47] Modulator 32.13 0.926 32.33 0.938 33.22 0.942 18.44 0.700 24.77 0.705
Mixture of LoRA Expert 32.28 0.930 32.62 0.945 33.39 0.949 20.59 0.719 25.05 0.716
w/o LoRA Expert 32.01 0.925 32.19 0.938 33.03 0.944 16.79 0.675 24.55 0.709
4.3 Ablation Study
We perform ablation studies to examine the role of each component in our proposed LoRA-IR. To comprehensively validate our method, we conduct experiments on the AllWeather [60] and mixed-degradation [91, 45] benchmarks. In Tab. 10, we start with LoRA-IR and systematically remove or replace modules, including the high-resolution techniques in DG-Router, the DAM module (we also attempt to use AdaLN [47] for feature modulation), and the mixture of LoRA expert design. We find that LoRA-IR consistently outperforms its ablated versions across all benchmarks, highlighting the critical importance of these components. Notably, our mixture of LoRA expert design significantly improves the model’s performance on mixed-degradation benchmarks, enhancing the model’s generalizability in real-world scenarios. More detailed ablations, model efficiency comparison and analysis are provided in the Appendix.
5 Conclusion
This paper introduces LoRA-IR, a flexible framework that dynamically leverages compact low-rank experts to facilitate efficient all-in-one image restoration. We propose a CLIP-based degradation-guided router (DG-Router) to extract robust degradation representations, requiring minimal training parameters and time. With the valuable guidance of the DG-Router, LoRA-IR dynamically integrates different low-rank experts, enhancing architectural flexibility while preserving computational efficiency. Across 14 image restoration tasks and 29 benchmarks, LoRA-IR demonstrates its state-of-the-art performance and strong generalizability.